write the standard question for each circle. then state the coordinates of its center and give its radius
x^2+y^2+4y=12

Answer by stanbon(75887)

write the standard equation for each circle. then state the coordinates of its center and give its radius
x^2+y^2+4y=12
------------------
(x-0)^2 + y^2 + 4y + 4 = 12 + 4
--------------------------------------
(x-0)^2 + (y+2)^2 = 16
----
Center:: (0,-2) ; radius = sqrt(16) = 4
---------------
Cheers,
Stan H.